OpenStudy (anonymous):

\[\sqrt{7} \div 2\sqrt{5} - \sqrt{7}\] rationalize the denominator.

OpenStudy (anonymous):

assuming this is \[\frac{\sqrt{7}}{2\sqrt{5}-\sqrt{7}}\] then multiply top and bottom by the conjugate of the denominator, in other words \[\frac{\sqrt{7}}{2\sqrt{5}-\sqrt{7}}\times \frac{2\sqrt{5}+\sqrt{7}}{2\sqrt{5}+\sqrt{7}}\]

OpenStudy (anonymous):

your denominator will then be \(a^2-b^2=20-7=13\) and your numerator will be whatever you get when you multiply out

OpenStudy (anonymous):

I don't get it...

OpenStudy (anonymous):

your second post

OpenStudy (anonymous):

which step?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

first of all it is always true that \((a+b)(a-b)=a^2-b^2\)

OpenStudy (anonymous):

so what does that make \((2\sqrt{5}-\sqrt{7})(2\sqrt{5}+\sqrt{7})\) ?

OpenStudy (anonymous):

it must be \[(2\sqrt{5})^2-\sqrt{7}^2\]

OpenStudy (anonymous):

when you square the first term you get \[2\sqrt{5}\times 2\sqrt{5}=4\times 5=20\] and when you square the second term you get \[\sqrt{7}\times \sqrt{7}=7\]

OpenStudy (anonymous):

and therefore \((2\sqrt{5}-\sqrt{7})(2\sqrt{5}+\sqrt{7})=(2\sqrt{5})^2-\sqrt{7}^2=20-7=13\)

OpenStudy (anonymous):

is that better?

OpenStudy (anonymous):

yeaah. thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you still have to multiply out in the numerator don't forget, but that is just the distributive law